OpenStudy (anonymous):

What is the inverse of the given relation? y = 3x + 9

OpenStudy (y2o2):

\[x = {{y-9} \over 3} = {1 \over3}y - 3\]

OpenStudy (alexwee123):

to find the inverse you would switch the x and y and then solve for y so.. y = 3x + 9 x=3y+9 3y=x-9 y=1/3*x-3

OpenStudy (anonymous):

Okay,, i had it right! Thank you

OpenStudy (y2o2):

you're welcome